OpenStudy (anonymous):

The perimeter of a rectangular field is 312 yards. If the length is 91 yards, what is its width?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Perimeter means the distance around the rectangle. $$\Huge Perimeter =2•length+2•width$$

OpenStudy (anonymous):

so..2*315?

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

Yes I am

OpenStudy (anonymous):



OpenStudy (anonymous):

so the perimeter is the length of the sides added up

OpenStudy (anonymous):

We have 2 sides which are 91 in length

OpenStudy (anonymous):

91+91=182

OpenStudy (anonymous):

So the width is 182?

OpenStudy (anonymous):

we are looking for 2 sides which have the same length

OpenStudy (anonymous):

no

OpenStudy (anonymous):

we will use x to represent the unknown width

OpenStudy (anonymous):

Since there are 2 of the sides with the same length for the width, we will call it 2x

OpenStudy (anonymous):

so the equation is 312 = 182 + 2x

OpenStudy (anonymous):

now just solve for x

OpenStudy (anonymous):

x=100?

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

nope, try and solve the equation

OpenStudy (anonymous):

what would be the first step?

OpenStudy (anonymous):

65?

OpenStudy (anonymous):

312=182+2x, what I did was I subtracted 182 from 312 and got 130 and 130/2 = 65

OpenStudy (anonymous):

that's correct, good job

OpenStudy (anonymous):

Will you help me with one more?

OpenStudy (anonymous):

open a new question and tag me

OpenStudy (anonymous):

okay, thank you!

OpenStudy (anonymous):

np